Find the general/particular solution of the following Differential Equations.

(Exact D.E)


[x Cos ( x + y ) + Sin ( x + y )] dy + x Cos (x+y) dy=0


1
Expert's answer
2021-09-09T09:59:45-0400
"(x \cos ( x + y ) + \sin ( x + y )) dx+ x\cos (x+y) dy=0"


"\dfrac{\partial P}{\partial y}=-x\sin(x+y)+\cos(x+y)"

"\dfrac{\partial Q}{\partial x}=\cos(x+y)-x\sin(x+y)"


"\dfrac{\partial P}{\partial y}=-x\sin(x+y)+\cos(x+y)=\dfrac{\partial Q}{\partial x}"

The differential equation 


"(x \cos ( x + y ) + \sin ( x + y )) dx+ x\cos (x+y) dy=0"

is an exact equation.

Then we write the system of two differential equations that define the function "u(x,y)"


"\begin{cases}
 \dfrac{\partial u}{\partial x}=x \cos ( x + y ) + \sin ( x + y ) \\
\\
\dfrac{\partial u}{\partial y}=x\cos(x+y)
\end{cases}"

Integrate the first equation over the variable "x"


"u(x, y)=\int(x \cos ( x + y ) + \sin ( x + y ))dx+\varphi(y)"

"\int x\cos(x+y)dx=x\sin(x+y)-\int\sin(x+y)dx"

"=x\sin(x+y)+\cos(x+y)+C_1"




"\int \sin(x+y)dx=-\cos(x+y)+C_2"

Then


"u(x, y)=\int(x \cos ( x + y ) + \sin ( x + y ))dx+\varphi(y)"


"=x\sin(x+y)+\cos(x+y)-\cos(x+y)+\varphi(y)"

"=x\sin(x+y)+\varphi(y)"

"u(x,y)=x\sin(x+y)+\varphi(y)"

Differentiate with respect to "y"


"\dfrac{\partial u}{\partial y}=x\cos(x+y)+\varphi'(y)=x\cos(x+y)"

"\varphi'(y)=0"

"\varphi(y)=-C"

The general solution of the exact differential equation is given by


"x\sin(x+y)=C"